BCorp needs 142,000 cartons of screws per year over the next 5 years. It'll cost 1,820,000 to install equipment. Depreciate the cost, straight-line to 0 over the projects life. In 5 years it can be salvaged at 152,000. Fixed production costs are 267,000 per year and variable production costs are $9.60 per carton. You need initial investment of NWC of $132,000. Tax rate is 22% and you require a 12% rate of return. Price per carton is $16.20.

A) What is the minimum number of cartons per year that can be supplied and still guarantee 0 NPV
Answer #1

CF0=-Cost of equipment-working capital
CF1, CF2, CF3, CF4, CF5=((number of cartons*(price per carton-variable cost per carton)-fixed costs per year-depreciation)*(1-tax rate)+depreciation)
Additional cash flow in year 5=Salvage value*(1-tax rate)+working capital
Depreciation=Initial cost of equipment/5
NPV=CF0+CF1/(1+r)+CF2/(1+r)^2+CF3/(1+r)^3+CF4/(1+r)^4+CF5/(1+r)^5+Additional cash flow in year 5/(1+r)^5

a. What is the minimum number of cartons per year that can be supplied and still break even?
At breakeven, NPV=0
-1820000+152000*(1-22%)/1.12^5-132000+132000/1.12^5+((Q*(16.20-9.60)-267000-1820000/5)*(1-22%)+1820000/5)/12%*(1-1/1.12^5)=0
=>Q=122424.866805